OpenStudy (turingtest):

Demonstrate that\[\frac{x^2+y^2}{4}\le e^{x+y-2}\]for all\[x\ge0,y\ge0\]

OpenStudy (turingtest):

A push in the right direction on this would suffice.

OpenStudy (turingtest):

...I hope.

OpenStudy (anonymous):

isn't right side circle?

OpenStudy (turingtest):

I suppose so. Isn't the left?

OpenStudy (anonymous):

yes, too early

OpenStudy (turingtest):

lol

OpenStudy (anonymous):

right side is \[\left(1+x+\frac{x^2}{2!}+\frac{x^3}{3!}+\text{...}\right)\left(1+y+\frac{y^2}{2!}+\frac{y^3}{3!}+\text{...}\right)\frac{1}{e^2}\]

OpenStudy (jamesj):

Write a function which is the difference, f(x,y) = e^(x+y-2) - (x^2+y^2)/4 Now, what I'd do is - Show that f(x,y) is non-negative on the axes (x,0) x ≥ 0, (0,y) y ≥ 0 - Show grad f points in the right direction; namely the x and y components are both positive, so we know f is increasing as we move away from those axes. - Then I'd find the right theorem so that I can formalize the conclusion of that second step and hence f(x,y) cannot be zero in that domain. If I get stuck on the last step, I might try and find a more elementary proof. I'll think about this a bit more.

OpenStudy (turingtest):

I'm pretty sure I will get stuck there, but thanks for the start, I'll give it a try.

OpenStudy (turingtest):

That seems equivalent to what James said. Only he used the difference of the functions.

OpenStudy (zarkon):

\[\frac{x^2+y^2}{4}=\frac{x^2+y^2}{4}=\frac{(x+y)^2-2xy}{4}\le\frac{(x+y)^2}{4}\]

OpenStudy (zarkon):

show \[\frac{z^2}{4}\le e^{z-2}\]...which is easy

OpenStudy (jamesj):

Yes, that's a more elementary proof.

OpenStudy (turingtest):

Induction?

OpenStudy (zarkon):

use calculus

OpenStudy (jamesj):

Write g(z) = e^(z-2) - z^2/4. Now use your elementary calculus to show that g(z) is positive for all z ≥ 0

OpenStudy (jamesj):

[ For the record, my method is intuitively sound, but I've been scribbling on some paper figuring out to formalize it. It's a bit complicated. One needs to a) establish the differential equations of gradient flow (d/dt)(x,y) = grad f; initial conditions f(x,0), f(0,y) [easy] b) Show that the solutions to those equations correspond to the values of f(x,y) for every point in the domain that is attained [not so bad] c) every point in the domain D = { (x,y) | x, y ≥ 0 } is attained by the gradient flow equations [tricky] In short, much better to use Zarkon's approach ;-) ]

OpenStudy (turingtest):

I figured as much, but I'm not used to doing proofs like this through calculus, so I'm rather of-balance. Easy he says...

OpenStudy (turingtest):

off*

OpenStudy (jamesj):

Actually, show that g(z) attains a minimum for z ≥0 and that minimum is positive.

OpenStudy (zarkon):

the min is actually zero (at z=2)...but that is ok

OpenStudy (jamesj):

non-negative, yes.

OpenStudy (turingtest):

But how in the heck do I solve\[e^{z-2}-\frac{z}{2}=0\]??? hmmm...

OpenStudy (turingtest):

It's obvious by inspection that you are right. z=2

OpenStudy (asnaseer):

there are actually two real solutions to that: http://www2.wolframalpha.com/input/?i=solve+e%5E%28x-2%29%3Dx%2F2

OpenStudy (asnaseer):

But I'm with you @TuringTest - how do we find the solution analytically here?

OpenStudy (turingtest):

Yeah, I did that as well, but how does that help if I can't find them? I don't even know what the product log function is, is that really necessary? Ah you'r stuck there too, at least I'm not alone :)

OpenStudy (asnaseer):

:)

OpenStudy (asnaseer):

I guess we don't have the photon drives installed in our brains - like @Zarkon and @JamesJ seem to have :-D

OpenStudy (turingtest):

yeah, I like how they say to use elementary calc to show that g(z) has a min at z=2 I know to take the derivative, no need to explain that. I don't know how to solve it is the problem. Here comes a tip...

OpenStudy (jamesj):

z = 2 is a solution of g'(z) = 0 and g'(z) > 0 for all z > 2 (why?), hence g(z) ≥0 for all z ≥ 2. Now, you'll need to be a bit creative (Rolle's Theorem? Mean Value Theorem? Change of sign of g''?), but stick at it for a bit and show there cannot be another zero between 0 and 2.

OpenStudy (turingtest):

That sounds promising :)

OpenStudy (asnaseer):

I went off on a tangent I proved this instead:\[\begin{align} f&=ex-e^x\\ f'&=e-e^x\\ \text{this is zero when }x&=1\\ f''&=-e^x\\ \text{this is negative at }x&=1\\ \text{therefore we have a maximum at }x&=1\\ f_{max}&=e-e=0\\ \therefore f\le0\\ \therefore ex-e^x&\le0\\ \therefore ex&\le e^x\\ \therefore x&\le e^{x-1}\\ \therefore xy&\le e^{x-1}e^{y-1}\\ \therefore xy&\le e^{x+y-2} \end{align}\]

OpenStudy (turingtest):

haha, still interesting though.

OpenStudy (turingtest):

Well I'm gonna try some of those tricks above...

OpenStudy (asnaseer):

yes - I'll give it a go as well...

OpenStudy (turingtest):

its easy enough to show that g'(z) > 0 for all z > 2 with the second derivative, but the rest... well let's see how creative you've been :)

OpenStudy (asnaseer):

\[\begin{align} g&=e^{z-2}-\frac{z^2}{4}\\ g'&=e^{z-2}-\frac{z}{2}\\ g''&=e^{z-2}-\frac{1}{2}\\ \end{align}\]we know z=2 gives g'=0. at z=2 g'' is positive so this represents a minimum \[\begin{align} g_{min}&=e^0-\frac{4}{4}=1-1=0\\ \therefore g&\ge0\\ \therefore e^{z-2} -\frac{z^2}{4}&\ge0\\ \therefore e^{z-2} &\ge \frac{z^2}{4} \end{align}\] oops - should have read your reply before carrying on :-)

OpenStudy (asnaseer):

I'd still like to know how to analytically find z=2 (and z=0.406376...) are the solutions to g'=0

OpenStudy (turingtest):

Hey that wasn't bad at all :P I guess I need to work on these, nice job! ^^^with you on that bit too.

OpenStudy (turingtest):

Although James made the argument that we need to prove that no other zeros exist between zero and two, did you do that?

OpenStudy (asnaseer):

no - I think we need to check the other solution as well, i.e. z=0.406376.. those are the only two that make g'=0

OpenStudy (turingtest):

Yeah, he says to show that there isn't a zero tin [0,2) but there is...

OpenStudy (turingtest):

in*

OpenStudy (jamesj):

a zero of g(x) I mean

OpenStudy (turingtest):

ah... I misunderstood.

OpenStudy (asnaseer):

at z=0.406376, g''=-0.296812 which means that represents a minimum

OpenStudy (jamesj):

a maximum

OpenStudy (asnaseer):

sorry - I meant max

OpenStudy (asnaseer):

therefore I think we have proved the assertion?

OpenStudy (turingtest):

But you had to use that solution that can't be found through inspection, how am I supposed to do that on the test? They don't let you bring Wolfram into the class.

OpenStudy (asnaseer):

yes - which is why I also am not 100% happy with the solution

OpenStudy (jamesj):

g(0) = e^-2. And we know that g(2) = 0. Further we know that for x just less than 2, g'(x) < 0. We also know that there is a zero of g(x), call it x1 and x1 < 0. Now, if there is another zero of g(x) between x1 and 2. Show that g(x) must have both a local max and local min in that range. But if that is the case, then g''(x) must have two zeros, but it doesn't.

OpenStudy (asnaseer):

I guess this is what @JamesJ is trying to say: